Mr. Wong is 4 times as old as his daughter Sylvia. In 10 years Mr. Wong will be 10 times as old as his daughter was 5 years ago. How old is each now?
Answer by macston(5194)

.
D=daughter's age; F=father's age=D+25
.
F+10yrs=2(D+10yrs) Substitute for F.
D+25yrs+10yrs=2D+20 yrs
D+35yrs=2D+20yrs
15yrs=D ANSWER 1: Daughter is 15 years old.
.
F=D+25yrs=15yrs+25yrs=40yrs ANSWER 2: Father is 40 years old
.
CHECK:
F+10yrs=2(D+10yrs)
40yrs+10yrs=2(15yrs+10yrs)
50yrs=2(26yrs)
50yrs=50yrs